The Madison Local High School marching band sold gift wrap to earn money for a band trip to Orlando,Florida .The gift wrap in solid colors sold for $4.00 per roll ,and the print gift wrap sold for $6.00 per roll .The total number of rolls sold was 480 ,and the total amount of money collected was $2340. How many rolls of each of each kind of gift wrap were sold?
Answer by scott8148(6628)

"The total number of rolls sold was 480" ___ s + p = 480 ___ -4s - 4p = -1920

"the total amount of money collected was $2340" ___ 4s + 6p = 2340

adding equations (to eliminate s) ___ 2p = 420

find p , then substitute back to find s